Find W1,W2,T1 and T2. Suppose that the pulleys are frictionless and the system is in equlibrium situation.

4001 50°
Answer #1

Twi & Wa T2 = Wz - Ti = W, - Ti 80140° + T2 6,50 =W2-③ W, 640 + Wz650 =W2 - 2 af from Ti Sess 40 = Tz sim so T, = Tzx 0.766 =